OpenStudy (anonymous):

i have a few please help What algebraic property does this statement show? If 7 = x, then x = 7. A. closure property B. associative property C. commutative property D. symmetric property

OpenStudy (ikram002p):

i guess C

OpenStudy (anonymous):

ok i have a few more

OpenStudy (ikram002p):

post them

OpenStudy (anonymous):

and thanks :)

OpenStudy (anonymous):

Solve. 3x – 5 = 7x – 17 A. 5 B. 3 C. 2 D. 6

OpenStudy (ikram002p):

3x – 5 = 7x – 17 -5x=-12 x=?

OpenStudy (anonymous):

your supposed to find what = the same hing so for example 3x5 – 5 = 7x5 – 17 and they have to = the same thing

OpenStudy (anonymous):

thing

OpenStudy (ikram002p):

wat !

OpenStudy (anonymous):

the first one =10 the second one = 18 they arnt equal so i move o the next one i figured this one out im still gonna need help after this one thoe

OpenStudy (anonymous):

its b

OpenStudy (ikram002p):

hmm i think all multiple choice u have are wrong see its linear equation so it wud be like this 3x – 5 = 7x – 17